It takes 7 hours for 8 cooks to prepare the food for a wedding rehearsal dinner. How long will it take 3 cooks to prepare the dinner? Also what kind variation (inverse or direct) applies to the situation?
Answer by stanbon(75887)

It takes 7 hours for 8 cooks to prepare the food for a wedding rehearsal dinner. How long will it take 3 cooks to prepare the dinner? Also what kind variation (inverse or direct) applies to the situation?
---
The more cooks there are, the less time it takes.
t = k/n where t is time and n is # of cooks.
-------------------------------
Inverse variation.
---
Solve for "k" using t=7 when n = 8:
7 =k/8
k = 56
----
Equation:
t = 56/n
------------
How long will it take 3 cooks to prepare the dinner?
t = 56/3
time = 56/3 = 18 2/3 hrs
==============================
Cheers,
Stan H.